The circumference of one wheel is 20mm longer than that of another wheel. The product of two circumference is 144300mm. Calculate the length of the two circumferences.

Grade:12th pass

1 Answers

Arun
5 years ago
given that c1=c2+20 
and 
c1*c2=144300 mm^2 
therefore 
(c2+20)*c2=144300 
 

hence c2^2+20*c2-144300=0 
this is a quadratic equation in c2 and can be solved using the standard 
procedure
as c2=(-20+760)/2=